### v1

# Reasoning Under Uncertainty: Exploring Probabilistic Reasoning Capabilities of LLMs

Abstract
Despite widespread success in language understanding and generation, large language models (LLMs) exhibit unclear and often inconsistent behavior when faced with tasks that require probabilistic reasoning. In this work, we present the first comprehensive study of the reasoning capabilities of LLMs over explicit discrete probability distributions. Given observations from a probability distribution, we evaluate models on three carefully designed tasks—mode identification, maximum likelihood estimation, and sample generation—by prompting them to provide responses to queries about either the joint distribution or its conditionals. These tasks thus probe a range of probabilistic skills, including frequency analysis, marginalization, and generative behavior. Through comprehensive empirical evaluations, we demonstrate that there exists a clear performance gap between smaller and larger models, with the latter demonstrating stronger inference and surprising capabilities in sample generation. Furthermore, our investigations reveal notable limitations, including sensitivity to variations in the notation utilized to represent probabilistic outcomes and performance degradation of over 60% as context length increases. Together, our results provide a detailed understanding of the probabilistic reasoning abilities of LLMs and identify key directions for future improvement.
1 Introduction
Large Language Models (LLMs) have achieved impressive performance across a wide range of Natural Language Processing (NLP) tasks, including question answering, summarization, and language understanding (Devlin et al., 2019; Raffel et al., 2020; Chowdhery et al., 2023; Joshi et al., 2017). Despite these advances, LLMs continue to exhibit notable limitations in probabilistic reasoning, as highlighted by several studies (Freedman and Toni, 2025; Gu et al., 2024; Ball et al., 2024). While there is growing interest in exploring their capabilities in probabilistic inference, it remains unclear how well these models can reason over explicit probability distributions. Existing research has primarily focused on logical reasoning, numerical problem-solving, and Bayesian inference (Chen et al., 2022; Ozturkler et al., 2022; Nafar et al., 2023, 2025; Qiu et al., 2025; Paruchuri et al., 2024), with limited attention given to reasoning over discrete probability distributions (Gu et al., 2024).
To address this gap, we present the first comprehensive evaluation of the reasoning capabilities of LLMs over discrete probability distributions. We develop a structured framework that prompts models with observations drawn from a probability distribution and evaluates their responses to queries targeting either the joint or conditional distributions. Our evaluation suite consists of three complementary tasks, mode identification, maximum likelihood estimation, and sample generation, that collectively probe distinct yet interconnected aspects of probabilistic reasoning. These tasks enable us to assess whether models can perceive the core characteristics of a distribution, infer underlying parameters that govern a distribution, and generate samples that reflect target distributions. By covering both joint and conditional settings, this study captures a broad range of reasoning abilities without requiring finetuning on probabilistic datasets.
Our experiments reveal clear distinctions in the capabilities of different models, with larger models and those distilled from them significantly outperforming the smaller ones in most settings. Across all three tasks, models consistently perform better on queries about joint distributions than on conditional ones, which require deeper understanding and reasoning. Surprisingly, some models demonstrate strong sampling capabilities, generating outputs that closely align with expected distributions. We also observe that model performance varies significantly depending on how outcomes of the distributions are represented in the prompts, highlighting pertinent biases that potentially arise due to the structure of the training data utilized. In addition, models exhibit significant weaknesses in counting, particularly as the context length increases, revealing a fundamental limitation in their ability to retain and reason over extended input sequences.
To summarize, we make the following contributions in this work:
  • We are the first to present an extensive evaluation of LLMs’ abilities to reason over discrete probability distributions across a variety of settings.
  • We introduce a diverse suite of complementary tasks— namely, mode identification, maximum likelihood estimation, and sampling—that investigate distinct aspects of probabilistic reasoning.
  • We conduct a systematic analysis across broad-ranging LLMs, revealing a substantial performance gap between small and large models, and uncover noteworthy sampling capabilities among certain models.
  • Furthermore, we highlight key limitations of current LLMs, including sensitivity to the notation of probabilistic outcomes, difficulty with accurate counting, and performance degradation when reasoning over long-context inputs.
2 Related Work
In-Context Learning: After the discovery of in-context learning (ICL) capabilities in large language models (LLMs), there has been a surge of interest in enhancing inference in zero-shot and few-shot settings. This paradigm enables models to perform tasks by conditioning on input examples without explicit parameter updates. Foundational studies have explored the mechanisms and effectiveness of ICL in various contexts (Brown et al., 2020; Min et al., 2022; Xie et al., 2021; Dai et al., 2022; Kojima et al., 2022; Wang et al., 2022). Building upon ICL, researchers have investigated methods to enable more advanced reasoning in LLMs, such as Chain-of-Thought (CoT) and Tree-of-Thoughts (ToT) frameworks that enhance problem-solving abilities (Yao et al., 2023; Wei et al., 2022).
Probabilistic Reasoning: Despite this advances, probabilistic reasoning and handling uncertainty remain challenging for LLMs (Nafar et al., 2023; Kadavath et al., 2022). Recent evaluations (Freedman and Toni, 2025) found that current models frequently violate basic probability rules such as complementarity and monotonicity, while (Ozturkler et al., 2022) proposed a structured approach, ThinkSum, decomposing probabilistic reasoning into distinct retrieval and aggregation stages. However, studies of (Gu et al., 2024) show that while LLMs understand probability concepts, they struggle to generate samples aligning with specified distributions. Another promising direction is Bayesian reasoning, where (Nafar et al., 2025) introduced the BLInD dataset to evaluate LLMs’ ability to perform Bayesian inference over interconnected variables. Furthermore, (Qiu et al., 2025) showed that training LLMs on predictions of an optimal Bayesian model significantly improves models’ capabilities.
Numerical Reasoning: Quantitative reasoning of LLMs is another aspect that has attracted notable attention. (Geva et al., 2020; Lewkowycz et al., 2022) demonstrated that numerical reasoning skills could be effectively introduced into LLMs through automated data generation and targeted training. (Chen et al., 2022) introduced the Program of Thoughts (PoT), where models express numerical reasoning as executable programs. Additionally, while (Paruchuri et al., 2024) showed that large models can reason about distributions when guided properly, (Liu et al., 2024) indicated their limited ability to infer probabilities from raw data or tables without special training and (Wang et al., 2025) highlighted biased and unreliable probability estimates by LLMs. Lastly, (Fu et al., 2024; Ball et al., 2024) revealed that LLMs struggle with counting tasks, and (Zhang et al., 2024) attributed these counting difficulties to architectural constraints and tokenization issues inherent to transformer models (Vaswani et al., 2017).
In summary, despite notable progress, key challenges remain in probabilistic reasoning, numerical accuracy, Bayesian inference, and counting, hence, addressing these gaps is essential for advancing the robustness and reliability of future LLMs.
3 Probabilistic Reasoning Framework
Figure 1: Prompt structure for evaluating LLMs on probabilistic reasoning tasks. Each prompt includes definitions of the random variables, observed samples or frequencies, a task specification (e.g., mode, MLE, sample generation), and output formatting instructions.
In this section, we introduce our framework for evaluating the probabilistic reasoning capabilities of LLMs. It begins with mathematical preliminaries, followed by details of the standardized prompt format we utilized. We then formalize three structured tasks, and describe our comprehensive evaluation pipeline. Together, these components enable a thorough and systematic comparison across different language models.
3.1 Preliminaries on Probabilistic Categorical Distributions
We formalize the probabilistic setup used in our study, wherein we assess whether LLMs can understand discrete probability distributions by evaluating their performance across several tasks.
Let $X=(X_{1},X_{2},\dots,X_{N})$ denote a tuple of $N$ categorical random variables, where each $X_{i}$ takes values in a finite label set $\mathcal{L}_{i}$. We refer to the elements of $\mathcal{L}_{i}$ as the possible labels of $X_{i}$. An outcome of the joint random variable $X$ is denoted by $\mathbf{x}=(x_{1},x_{2},\dots,x_{N})$, where $x_{i}\in\mathcal{L}_{i}$ for all $i\in[N]$. Let $\mathcal{L}=\mathcal{L}_{1}\times\mathcal{L}_{2}\times\dots\times\mathcal{L}_{N}$ denote the set of all possible joint outcomes. The size of the joint outcome space is then given by
$|\mathcal{L}|=\prod_{i=1}^{N}|\mathcal{L}_{i}|.$ (1)
To grant the LLM access to a distribution, we provide a set of $K$ independent observations drawn from the joint distribution. Let $\mathcal{O}=\{\mathbf{x}^{(1)},\mathbf{x}^{(2)},\dots,\mathbf{x}^{(K)}\}$ denote the observed samples. This information is incorporated into the LLM’s context using one of two following formats: (1) by providing the empirical frequency of each possible outcome $\mathbf{x}$, or (2) by directly including the set of observations $\mathcal{O}$ in the prompt.
3.2 Prompt Format
As shown in Figure 1, each prompt begins with a fixed segment that defines the random variables $X_{1},X_{2},\dots,X_{N}$, their corresponding label sets $\mathcal{L}_{1},\mathcal{L}_{2},\dots,\mathcal{L}_{N}$ and the joint probability distribution $P(X_{1},X_{2},\dots,X_{N})$. This is followed by observations from the joint distribution, provided either as empirical frequencies or raw samples. The prompt concludes with a task-specific query and clear instructions, explicitly discouraging code generation and specifying the expected output format. A complete list of prompts used for each task and setting is provided in the Appendix A.
3.3 Task Definition and Performance Metrics
In this section, we describe our evaluation suite and the corresponding metrics used to assess the capability of LLMs in understanding categorical probability distributions. For each of the following tasks, we consider two settings: one involving queries about joint distributions, and the other involving queries about conditional distributions.
Mode Identification: Given a set of $K$ observations $\mathcal{O}=\{\mathbf{x}^{(1)},\dots,\mathbf{x}^{(K)}\}$ provided to the model, this task evaluates whether the model can identify the mode of the empirical distribution.
Joint task. We measure if the model can find the mode of the joint distribution i.e.,
$\hat{\mathbf{x}}_{\text{mode}}=\operatorname*{arg\,max}_{\mathbf{x}\in\mathcal{L}}\frac{1}{K}\sum_{k=1}^{K}\mathds{1}{[\mathbf{x}^{(k)}=\mathbf{x}]},$
Conditional task. We condition on a random variable $X_{c}$ and evaluate the mode of a query variable $X_{q}$. Given a conditioning value $X_{c}=x_{c}$ for some $x_{c}\in\mathcal{L}_{c}$, we define:
$\hat{x}_{q}=\operatorname*{arg\,max}_{x_{q}\in\mathcal{L}_{q}}\hat{p}(X_{q}=x_{q}\mid X_{c}=x_{c}),$
where $\hat{p}$ is the empirical conditional distribution estimated from the subset of $\mathcal{O}$ satisfying $X_{c}=x_{c}$. We report accuracy for this task, indicating whether the model correctly identifies the most likely outcome.
Maximum Likelihood Estimation (MLE): Given the observed dataset $\mathcal{O}$, the model is asked to estimate the empirical distribution over possible outcomes—either for the joint distribution or in a conditional manner.
Joint task: The model estimates $\hat{p}(X=\mathbf{x})$ for all $\mathbf{x}\in\mathcal{L}$. We measure the total variation distance (TVD) between the predicted empirical distribution $\hat{p}$ and the true empirical distribution $p$ from $\mathcal{O}$, i.e.,
$\frac{1}{2}\sum_{\mathbf{x}\in\mathcal{L}}\big{\lvert}p(X=\mathbf{x})-\hat{p}(X=\mathbf{x})\big{\rvert}$.
Conditional task: Given $X_{c}=x_{c}$, the model estimates $\hat{p}(X_{q}=x_{q}\mid X_{c}=x_{c})$ for all $x_{q}\in\mathcal{L}_{q}$. We once again report the total variation distance for this task, i.e.,
$\frac{1}{2}\sum_{x_{q}\in\mathcal{L}_{q}}\big{\lvert}p(X_{q}=x_{q}\mid X_{c}=x_{c})-\hat{p}(X_{q}=x_{q}\mid X_{c}=x_{c})\big{\rvert}$.
Sample Generation: In this task, the model is prompted to generate new samples from the joint or conditional distribution based on the provided observations $\mathcal{O}$. Unlike the previous tasks focused on estimation or selection, this task evaluates the model’s generative capability—i.e., whether it can reproduce the empirical distribution through sampling.
Joint task: The model is prompted to generate $K^{\prime}$ samples based on the joint distribution provided in the context. Let $\hat{\mathcal{O}}$ be the samples generated by the LLM, and let $q$ denote their empirical frequency distribution. We report the total variation distance, which is defined in this setting as:
$\frac{1}{2}\sum_{\mathbf{x}\in\mathcal{L}}\big{\lvert}p(X=\mathbf{x})-q(X=\mathbf{x})\big{\rvert}$.
Conditional task: The model is prompted to generate $K^{\prime}$ samples of the random variable $X_{q}$ conditioned upon $X_{c}=x_{c}$. Let $q(X_{q}\mid X_{c}=x_{c})$ be the empirical frequency distribution over the samples generated by the LLM. We report the total variation distance, which is defined in this setting as:
$\frac{1}{2}\sum_{x_{q}\in\mathcal{L}_{q}}\big{\lvert}p(X_{q}=x_{q}\mid X_{c}=x_{c})-q(X_{q}=x_{q}\mid X_{c}=x_{c})\big{\rvert}$
By evaluating models on joint and conditional distributions, we examine their ability to handle probabilistic relationships in various settings. Each task focuses on a specific aspect, providing a comprehensive assessment of how well LLMs understand and process discrete probability distributions.
3.4 Evaluation Setup
As outlined in Section 3.2, output formats are explicitly defined in the prompts to guide models toward consistent responses. Through extensive prompt engineering, task-specific formats are identified that models generally find easier to follow. When responses match the expected format, final answers are extracted using regular expressions, and the corresponding evaluation metrics are applied as mentioned above. While larger models typically follow the instructions closely, smaller models often return answers in inconsistent or alternative formats. To handle such cases, a capable model, Llama3.3-70B, is used as an LLM judge. It receives the task description, the expected answer, and the model’s response, and determines whether the output reflects the correct answer. This approach supports a robust evaluation and helps ensure the reliability of reported results. We provide additional details regarding the usage of the LLM Judge, in the Appendix C due to paucity of space.
4 Robustness Analysis
In this section, we investigate pertinent factors that influence LLMs’ performance on probabilistic reasoning tasks by focusing on three key aspects: (1) how performance changes as the task difficulty increases, (2) how sensitive models are to variations in the label notation used to represent outcomes of the random variables, and (3) how model behavior differs when input prompts include direct, raw sample observations instead of empirical frequencies. These aspects expose distinct challenges for models and offer deeper insights into their strengths and weaknesses.
4.1 Scaling Complexity
To examine how task complexity affects model performance, we vary the size of the outcome or label space $|\mathcal{L}|$ for the joint distributions considered, which directly influences the difficulty of all three tasks. As shown in Equation (1) increasing either the number of random variables $N$ or the number of labels $|\mathcal{L}_{i}|$ for each variable causes exponential or multiplicative growth of the joint space respectively, requiring models to reason over a larger set of outcomes and their associated probabilities. In the mode identification task, this expansion increases the number of candidates the model must compare, making it harder to identify the most probable outcome. In the MLE task, a larger joint space implies that the model must estimate a greater number of probabilities, making accurate prediction more challenging. Similarly, in the sampling task, the model is expected to generate samples that reflect a more complex distributed set of outcomes, demanding stronger generalization and a better understanding of the underlying probabilities.
4.2 Label Sensitivity
Our experiments reveal that LLMs are highly sensitive to the choice of labels utilized in the set $\mathcal{L}_{i}$, with performance varying significantly across different label categories. Interestingly, even when the set of labels and frequency counts are fixed, simply changing which outcome is associated with which frequency can lead to noticeable shifts in model performance. This behavior likely stems from biases in the training data and the tendency of the model to interpret labels as part of the linguistic context, rather than discrete outcomes in a probabilistic space. Additionally, sensitivity to label variation differs across models, with smaller models generally showing greater variability. To quantify this effect, we design a set of experiments to measure models’ robustness to superficial changes in label notation.
4.3 Estimations using Samples-in-Context Instead of Frequencies
In this setting, we explore the effect of including the raw observations $\{\mathbf{x}^{(1)},\mathbf{x}^{(2)},\dots,\mathbf{x}^{(K)}\}$ in the input prompts, rather than their empirical frequencies. By doing so, models should count the frequencies of each outcome and solve the probabilistic tasks accordingly. This change shifts the problem from simple frequency analysis to counting and subsequent processing of the counts so obtained. The longer the context, the more challenging it becomes for the model to correctly interpret the frequency of occurrences, leading to a sharp decline in performance. In the next section, we illustrate a comparison between these two approaches and highlight the weaknesses of powerful models in counting and managing a large number of samples.
5 Experiments
This section outlines the experimental setup used to evaluate LLMs on probabilistic reasoning tasks. It introduces the models selected for evaluation and describes key implementation details. Results are presented for the three core tasks, followed by additional studies that examine factors influencing model performance.
5.1 Model Selection
In this study, we evaluate several instruction-finetuned LLMs, selected to represent a wide range of model sizes and capabilities. The models evaluated include Llama3.1-8B, Llama3.3-70B (Grattafiori et al., 2024), Qwen2.5-7B (Yang et al., 2024), DeepSeek-R1-Distill-Qwen-7B (Guo et al., 2025), GPT-4o-mini (Hurst et al., 2024), and OpenAI’s recently published model, GPT-4.1-mini (Achiam et al., 2023). These models span a broad spectrum of parameter sizes, from 7B to 70B, including specialized reasoning models, offering a comprehensive comparison of how different architectures handle various settings.
Table 1: Accuracy comparison on mode identification tasks across different joint sizes $|\mathcal{L}|$. cond-mode requires reasoning over conditional distributions, making it generally more challenging. While larger models—or those distilled from them—maintain robust performance across all settings, smaller models show a clear decline in accuracy as $|\mathcal{L}|$ increases.
Model joint-mode cond-mode $|\mathcal{L}|=12$ $|\mathcal{L}|=27$ $|\mathcal{L}|=36$ $|\mathcal{L}|=54$ $|\mathcal{L}|=12$ $|\mathcal{L}|=27$ $|\mathcal{L}|=36$ $|\mathcal{L}|=54$ Llama3.1-8B 0.59 0.57 0.35 0.21 0.83 0.79 0.68 0.62 Qwen2.5-7B 0.93 0.83 0.81 0.65 0.69 0.67 0.63 0.60 DeepSeek-R1-Distill-Qwen-7B 0.95 0.93 0.91 0.86 0.93 0.93 0.91 0.83 Llama3.3-70B 1.00 1.00 1.00 0.97 1.00 0.99 0.96 0.95 GPT-4o-mini 0.99 0.98 0.98 0.80 0.99 0.96 0.95 0.88 GPT-4.1-mini 1.00 1.00 0.99 0.96 1.00 1.00 1.00 1.00
5.2 Experimental Setup
In our experiments, we follow the standard naming convention for random variables, using names such as $X,Y,Z,T,$ etc., and set $K:=5\times|\mathcal{L}|$, meaning the number of observations is five times larger than the joint space. We then select $|\mathcal{L}|$ values that sum to $K$, as frequencies, and randomly assign them to different joint outcomes. However, as discussed in Section  4.2, models are sensitive to how these frequencies are linked to specific outcomes, even when both the set of labels and frequency values are unchanged. To reduce the impact of this sensitivity, we create ten different ways of assigning frequencies to joint outcomes and generate ten prompts for each assignment, resulting in 100 prompts per task. This helps ensure that variations in outcome-frequency mapping minimally impacts the evaluation results. For larger $|\mathcal{L}|$, we carefully select the frequency values to generate a distribution with entropy comparable to that of the simpler distributions, ensuring consistency across experiments. Given the models’ constraints in handling raw samples in the input, we incorporate the empirical frequency of observed samples in most experiments and analyze the comparison between the two approaches in Section 5.4.2.
Table 2: Average TVD across 100 prompts for maximum likelihood estimation tasks at varying joint sizes $|\mathcal{L}|$. joint-MLE is consistently easier with near-zero TVD, while cond-MLE seems more challenging and highlights a clear gap between models.
Model joint-MLE cond-MLE $|\mathcal{L}|=12$ $|\mathcal{L}|=27$ $|\mathcal{L}|=36$ $|\mathcal{L}|=54$ $|\mathcal{L}|=12$ $|\mathcal{L}|=27$ $|\mathcal{L}|=36$ $|\mathcal{L}|=54$ Llama3.1-8B 1e-05 0.001 0.003 0.02 0.089 0.115 0.151 0.198 Qwen2.5-7B 0 2e-05 4e-04 0.006 0.077 0.106 0.140 0.177 DeepSeek-R1-Distill-Qwen-7B 0 5e-04 0.001 0.02 0.038 0.048 0.069 0.100 Llama3.3-70B 0 0 0 4e-04 0.009 0.003 0.017 0.084 GPT-4o-mini 0 2e-05 7e-04 1e-04 0 0.008 0.017 0.084 GPT-4.1-mini 0 0 1e-05 1e-04 0 5e-06 5e-05 0.016
5.3 Task-Specific Experimental Results
This section presents a detailed analysis of the experimental results for each task. To evaluate performance on conditional queries, one random variable $X_{c}$ and one of its possible outcomes $x_{c}$ are randomly selected as evidence. Another variable is then chosen as the query variable $X_{q}$ and the model is asked to solve one of the tasks based on the conditional distribution $P(X_{q}\mid X_{c}=x_{c})$. For all tasks, each model is prompted 100 times using 10 different prompts, and the outputs are parsed to extract the final answer for each query.
Mode Identification: In this task, models are asked to identify the mode of either the joint distribution (joint-mode) or a conditional distribution (cond-mode). The joint-mode task requires selecting the outcome with the highest overall frequency, without additional computation. In contrast, the cond-mode task involves identifying the most frequent outcome given a specified condition, requiring models to filter and reason over a subset of the distribution. Table 1 reports model accuracy as the joint space size $|\mathcal{L}|$ increases. Although the cond-mode task involves fewer possible outputs, the added conditional reasoning makes it more challenging. Consequently, most models achieve slightly lower accuracy on cond-mode compared to joint-mode. While smaller models struggle as the distribution grows, Llama3.3-70B and GPT4.1-mini maintain near-perfect accuracy across all joint sizes. The details of different distributions can be found in Appendix B.
Maximum Likelihood Estimation: In this task, models are asked to predict either the probability of all joint outcomes (joint-MLE) or the conditional probabilities $P(X_{q}=x_{q}\mid X_{c}=x_{c})$ for each $x_{q}\in\mathcal{L}_{q}$ (cond-MLE). For evaluation, the TVD is computed for each prompt, and the average is reported over 100 prompts. As shown in Table 2, the joint-MLE task appears to be straightforward for all models, with even smaller models achieving near-zero TVD across all joint sizes. In contrast, the cond-MLE task is more challenging, as it requires models to derive conditional probabilities from the joint distribution. This added complexity results in consistently higher TVD values, with a clear performance gap between smaller and larger models.
Sample Generation: This task evaluates the generative abilities of LLMs by requiring them to generate $K^{\prime}$ samples autoregressively from either the joint distribution (joint-sampling) or a conditional distribution (cond-sampling). We set $K^{\prime}:=7\times|\mathcal{L}|$ to ensure that the generated samples are sufficient to approximate the target distribution. Since it is inherently difficult to reproduce a target empirical distribution with a finite number of samples, we also compute the TVD between samples drawn from Python’s random library and the true distribution. This serves as a baseline for comparing the quality of samples generated by the models.
As shown in Figure 2, TVD in the joint-sampling task increases with $|\mathcal{L}|$. Interestingly, Llama3.3-70B and GPT4.1-mini achieve TVD values even lower than those from Python’s random sampler, which includes sampling noise, suggesting that the underlying distribution of their generated samples is remarkably close to the target. In contrast, all models struggle with the cond-sampling task, as it implicitly requires the models to first estimate the conditional probabilities and then generate samples that align with the resulting distribution. Surprisingly, even though no model reaches a satisfactory level of performance, DeepSeek-R1-Distill-Qwen, designed for reasoning tasks, achieves the best performance on this task, outperforming even the strongest models in other settings.
Beyond distributional evaluation, we also assessed whether model-generated samples exhibit true independence. Using the Durbin–Watson statistic (Durbin and Watson, 1951) and transition matrix entropy (Shannon, 1948), we found evidence of correlations in all models’ outputs, indicating that generations are not fully independent and are influenced by previously produced samples, consistent with the autoregressive nature of LLMs. In a related experiment, GPT-4.1-mini was prompted to generate one sample at a time across multiple prompts, we observed that it predominantly produced high-frequency outcomes while almost entirely avoiding rare ones. These findings highlight that current LLMs not only struggle with conditional sampling but also face challenges in generating truly independent samples. Full experimental details and results are provided in Appendix D.
To validate our findings on real-world data, we evaluated all six task configurations using the Mushroom dataset (mus, 1981). We focused on a subset of this dataset, resulting in a highly skewed distribution. The detailed results for each task are provided in Appendix E, enabling a comprehensive comparison of model performance under realistic conditions.
Figure 2: Average TVD across 100 prompts for joint-sampling (left) and cond-sampling (right) tasks, with models queried to generate $7\times|\mathcal{L}|$ samples. Here, the sampling baseline is computed as the TVD between $K^{\prime}$ samples generated by Python’s random function and the true distribution.
5.4 Robustness Analysis Results
This section presents the results of analyzing two key aspects: the sensitivity of models to variations in outcomes of the distributions, and the effect of providing raw samples in the prompt instead of their empirical frequencies.
5.4.1 Label Sensitivity Results
To investigate the impact of label choice on model performance, a targeted set of experiments is conducted on three models: the two top-performing models, Llama3.3-70B and GPT-4.1-mini, and a smaller model, Qwen2.5-7B. All three are evaluated on the joint-mode task, which they handle reliably under standard conditions, using joint space size of $|\mathcal{L}|=54$ and ten distinct label categories. These categories span a broad range of semantics, including computer science terms, human names, US states, countries, fruits, universities, majors, car brands, neurology terms, and food. By systematically varying the labels used to represent the outcomes of the random variables, the goal is to assess how surface-level changes in label representation influence model behavior.
Figure 3 (right) presents the accuracy of each model across the different label categories. GPT-4.1-mini shows strong consistency, with minimal variation in accuracy across categories, indicating a high level of robustness to changes in label sets. Llama3.3-70B performs well overall but exhibits moderate sensitivity, with performance varying more noticeably across label sets. In contrast, Qwen2.5-7B is highly sensitive to label changes, suffering substantial drops in accuracy depending on the category. These findings suggest that while larger models or those distilled from them can generalize well across prompt variations, smaller models are significantly vulnerable to the way information is presented, even when the underlying task remains unchanged.
5.4.2 Samples-in-Context Results
Next, we investigate how model performance is affected when raw observed samples are provided in the prompt instead of their empirical frequencies. This setting is evaluated using the two best-performing models, Llama3.3-70B and GPT-4.1-mini, as smaller models consistently fail under these conditions. We repeat the joint-mode task across different joint space sizes, but rather than supplying frequency counts for each outcome, we provide the raw samples directly as mentioned. This forces the models to perform implicit counting—first identifying the frequency of each outcome from the raw data, and then determining the mode based on that.
The results reveal a sharp decline in performance: both models, which had previously achieved near-perfect accuracy with explicit frequencies, struggle significantly in this setting. As the joint space size $|\mathcal{L}|$ increases and the input grows longer, their ability to track counts and identify the correct outcome diminishes rapidly. This performance drop is illustrated in Figure 3 (left), which shows a near-linear decline in accuracy with increasing joint size. The findings point to a core limitation: the models’ inability to reliably perform counting over long contexts. Among the two, GPT-4.1-mini consistently outperforms Llama3.3-70B, indicating greater robustness to the added complexity and input length. Further evidence from Appendix F shows that even when the joint space size is fixed at $|\mathcal{L}|=12$, increasing the number of observed samples $K$ results in additional accuracy degradation. This suggests that the challenge is not only tied to task complexity but also to the models’ limited capacity for reasoning over long sequences.
To explore potential mitigation strategies, we conducted an additional experiment with GPT-4.1-mini where the model was given access to a code interpreter. In this setup, the model was encouraged to write Python code to count frequencies before solving the task. This hybrid approach substantially improved performance, nearly matching the results from the frequency-provided setting. The detailed results of these experiments are presented in Appendix G.
Figure 3: (Left) Accuracy comparison on the joint-mode task when models are given either raw samples or frequency counts. Accuracy sharply declines with increasing joint space size when raw samples are used, revealing the difficulty of reasoning over longer contexts without explicit frequency information. GPT-4.1-mini shows greater robustness than Llama3.3-70B in this setting. (Right) Accuracy of models across ten different label categories in the joint-mode task. GPT-4.1-mini remains consistently robust across categories and Qwen2.5-7B exhibits significant sensitivity.
6 Conclusion
This work provides the first comprehensive evaluation of large language models’ abilities to reason over discrete probability distributions. By examining a range of structured tasks including mode identification, maximum likelihood estimation, and sample generation, we observe clear differences in capability between models. Larger models or those distilled from them generally perform well across tasks, but their strengths are especially evident in sample generation, where the produced samples align more closely with the target distribution than those of a true random sampler. However, we find that performance consistently degrades when tasks are framed as conditional rather than joint distributions, revealing an ongoing challenge in conditional reasoning. Moreover, models remain sensitive to superficial changes in outcome labels and consistently fail when required to infer probabilities from raw samples instead of provided frequencies. Together, these findings highlight both the significant progress of modern LLMs in probabilistic reasoning and the fundamental challenges that must be addressed to achieve reliable reasoning under uncertainty.
Limitations
This study can extend in many directions, each requiring substantial time, effort, and computational resources. To maintain focus and feasibility, we limited our scope to a subset of these directions and aimed to investigate them thoroughly. One key limitation is the scale of the joint distributions used in our experiments. We limit the joint space size at 54 since smaller models struggled with longer contexts, leading to unstable and unreliable performance. Scaling to larger joint spaces would further degrade performance and reduce comparability across model sizes. Another limitation lies in our choice of distribution types. While analyzing both discrete and continuous distributions would provide a more complete picture of LLMs’ capabilities, we deliberately focused on discrete distributions. We identified a significant gap in understanding LLMs’ reasoning over discrete distributions, and we believe this gap deserves focused attention as a foundational problem that presents distinct challenges from continuous distributions. Finally, we restricted our exploration to a limited subset of distributions due to the substantial effort and time required to evaluate a broader range. Future work is needed to explore and uncover these remaining directions.
Acknowledgements
This project was supported in part by a grant from an NSF CAREER AWARD 1942230, the ONR PECASE grant N00014-25-1-2378, ARO’s Early Career Program Award 310902-00001, Army Grant No. W911NF2120076, the NSF award CCF2212458, NSF Award No. 2229885 (NSF Institute for Trustworthy AI in Law and Society, TRAILS), a MURI grant 14262683, DARPA AIQ DARPA AIQ grant HR00112590066 and an award from meta 314593-00001.
Appendix
This section includes complete prompt examples for each task, the exact frequency distributions used for various joint space sizes, the prompt template used to evaluate outputs with the LLM judge and supplemental results for samples-in-Context settings.
Appendix A Task Prompts
Each task is evaluated with two prompt variants: one for reasoning over joint distributions and one for conditional distributions. We designed three structured tasks to evaluate different aspects of probabilistic reasoning. Below are the representative prompts used for each task.
A.1 Mode Identification
Below, is an example of a prompt used for joint-mode task for $|\mathcal{L}|=12$:
Consider 3 discrete random variables X, Y, Z, where X can take outcomes from A, B, Y can take outcomes from C, D, Z can take outcomes from E, F, G. In 60 independent samples drawn from their joint distribution P(X, Y, Z), the observed frequencies are:
- (A, C, E): 5
- (A, C, F): 7
- (A, C, G): 3
- (A, D, E): 10
- (A, D, F): 2
- (A, D, G): 3
- (B, C, E): 6
- (B, C, F): 8
- (B, C, G): 2
- (B, D, E): 3
- (B, D, F): 3
- (B, D, G): 8
Task: Identify the mode (most probable outcome) of the joint distribution P(X, Y, Z).
Instructions:
1. Do not write any code or pseudocode.
2. Strictly follow the output format below.
3. You may explain your reasoning, but the final answer should be explicitly summarized at the end. You get negative penalty for not following the output format.
Output Format that you should strictly follow:
Mode = (, , )
Below is a variation of the above prompt that requires solving the task on a conditional distribution (cond-mode).
Consider 3 discrete random variables X, Y, Z, where X can take outcomes from A, B, Y can take outcomes from C, D, Z can take outcomes from E, F, G. In 60 independent samples drawn from their joint distribution P(X, Y, Z), the observed frequencies are:
- (A, C, E): 5
- (A, C, F): 7
- (A, C, G): 3
- (A, D, E): 10
- (A, D, F): 2
- (A, D, G): 3
- (B, C, E): 6
- (B, C, F): 8
- (B, C, G): 2
- (B, D, E): 3
- (B, D, F): 3
- (B, D, G): 8
Task: Identify the mode (most probable outcome) of the conditional distribution P(Z | X = A).
Instructions:
1. Do not write any code or pseudocode.
2. Strictly follow the output format below.
3. You may explain your reasoning, but the final answer should be explicitly summarized at the end. You get negative penalty for not following the output format.
Output Format that you should strictly follow:
Most probable value of Z given X=A is
A.2 Maximum Likelihood Estimation
The model is asked to estimate the probabilities of the joint distributions using maximum likelihood estimates from observed frequencies.
A prompt sample for joint-MLE task for $|\mathcal{L}|=12$:
Consider 3 discrete random variables X, Y, Z, where X can take outcomes from A, B, Y can take outcomes from C, D, Z can take outcomes from E, F, G. In 60 independent samples drawn from their joint distribution P(X, Y, Z), the observed frequencies are:
- (A, C, E): 5
- (A, C, F): 7
- (A, C, G): 3
- (A, D, E): 10
- (A, D, F): 2
- (A, D, G): 3
- (B, C, E): 6
- (B, C, F): 8
- (B, C, G): 2
- (B, D, E): 3
- (B, D, F): 3
- (B, D, G): 8
Task: Predict the maximum likelihood estimation (MLE) of the joint probability distribution P(X, Y, Z) based on these 60 samples.
Instructions:
1. Think step by step and solve this mathematically using probability theory - do not write any code or pseudocodes as you get negative penalty for that.
2. Clearly state the final estimated probabilities for each (X, Y, Z) outcome. Output probabilities should be expressed as float numbers with up to four decimal points.
3. You may explain your reasoning, but the final answer should be explicitly summarized at the end. only your final answer will be graded.
Output Format:
- Final probabilities as float numbers should be listed as:
P(A, C, E) = [value]
P(A, C, F) = [value]
P(A, C, G) = [value]
P(A, D, E) = [value]
P(B, D, F) = [value]
P(B, D, G) = [value]
Below is a variation of the above prompt that requires solving the task on a conditional distribution (cond-MLE).
Consider 3 discrete random variables X, Y, Z, where X can take outcomes from A, B, Y can take outcomes from C, D, Z can take outcomes from E, F, G. In 60 independent samples drawn from their joint distribution P(X, Y, Z), the observed frequencies are:
- (A, C, E): 5
- (A, C, F): 7
- (A, C, G): 3
- (A, D, E): 10
- (A, D, F): 2
- (A, D, G): 3
- (B, C, E): 6
- (B, C, F): 8
- (B, C, G): 2
- (B, D, E): 3
- (B, D, F): 3
- (B, D, G): 8
Task: Predict the maximum likelihood estimation (MLE) of the conditional distribution P(X | Z = E).
Instructions:
1. Think step by step and solve this mathematically using probability theory - do not write any code or pseudocodes as you get negative penalty for that.
2. Clearly state the final estimated probabilities as float numbers with up to four decimal points.
3. You may explain your reasoning, but the final answer should be explicitly summarized at the end.
Output Format:
- Final probabilities as float numbers should be listed as:
P(X=A | Z=E) = [value]
P(X=B | Z=E) = [value]
A.3 Sample Generation
In this task, the model is asked to generate samples either from the joint distribution or from a conditional distribution, and we compare the empirical distribution of generated samples with the true distribution.
A prompt sample for joint-sampling task for $|\mathcal{L}|=12$:
Consider 3 discrete random variables X, Y, Z, where X can take outcomes from A, B, Y can take outcomes from C, D, Z can take outcomes from E, F, G. In 60 independent samples drawn from their joint distribution P(X, Y, Z), the observed frequencies are:
- (A, C, E): 5
- (A, C, F): 7
- (A, C, G): 3
- (A, D, E): 10
- (A, D, F): 2
- (A, D, G): 3
- (B, C, E): 6
- (B, C, F): 8
- (B, C, G): 2
- (B, D, E): 3
- (B, D, F): 3
- (B, D, G): 8
Task: Generate EXACTLY 84 random samples from the joint distribution P(X, Y, Z).
Instructions:
1. OUTPUT MUST BEGIN with ### Output on a new line.
2. List EXACTLY 84 samples numbered 1 to 84.
3. Do not write any code or pseudocode.
4. Strictly follow the output format below.
5. You will be penalized for writing codes or not following the output format.
Output Format that you should strictly follow:
### Output
1. (X, Y, Z)
2. (X, Y, Z)
84. (X, Y, Z)
Where X can take outcomes from A, B, Y can take outcomes from C, D, Z can take outcomes from E, F, G.
Below is a variation of the above prompt that requires solving the task on a conditional distribution (cond-sampling).
Consider 3 discrete random variables X, Y, Z, where X can take outcomes from A, B, Y can take outcomes from C, D, Z can take outcomes from E, F, G. In 60 independent samples drawn from their joint distribution P(X, Y, Z), the observed frequencies are:
- (A, C, E): 5
- (A, C, F): 7
- (A, C, G): 3
- (A, D, E): 10
- (A, D, F): 2
- (A, D, G): 3
- (B, C, E): 6
- (B, C, F): 8
- (B, C, G): 2
- (B, D, E): 3
- (B, D, F): 3
- (B, D, G): 8
Task: Generate EXACTLY 84 random samples from the conditional distribution P(Z | X=A).
Instructions:
1. OUTPUT MUST BEGIN with ### Output on a new line.
2. List EXACTLY 84 samples numbered 1 to 84.
3. Do not write any code or pseudocodes.
4. Strictly follow the output format below.
5. You will be penalized for writing codes or not following the output format.
Output Format that you should strictly follow:
### Output
1. Z
2. Z
84. Z
Where Z can take outcomes from E, G, F.
Appendix B Distributions Used for Each Joint Space Size
Distribution used for $|\mathcal{L}|=4$
("A", "C"): 6, ("A", "D"): 4, ("B", "C"): 8, ("B", "D"): 2
Distribution used for $|\mathcal{L}|=6$
("A", "C"): 7, ("A", "D"): 2, ("A", "E"): 4, ("B", "C"): 9, ("B", "D"): 2, ("B", "E"): 6
Distribution used for $|\mathcal{L}|=8$
("A", "C", "E"): 7, ("A", "C", "F"): 4, ("A", "D", "E"): 9,
("A", "D", "F"): 2, ("B", "C", "E"): 4, ("B", "C", "F"): 3,
("B", "D", "E"): 6, ("B", "D", "F"): 5
Distribution used for $|\mathcal{L}|=9$
("A", "D"): 6, ("A" , "E"): 3, ("A", "F"): 4, ("B", "D"): 9, ("B", "E"): 2, ("B", "F"): 5, ("C", "D"): 4, ("C", "E"): 3, ("C", "F"): 4
Distribution used for $|\mathcal{L}|=12$
("A", "C", "E"): 8, ("A", "C", "F"): 2, ("A", "C", "G"): 7, ("A", "D", "E"): 3, ("A", "D", "F"): 3, ("A", "D", "G"): 3, ("B", "C", "E"): 6, ("B", "C", "F"): 10, ("B", "C", "G"): 2, ("B", "D", "E"): 8, ("B", "D", "F"): 5, ("B", "D", "G"): 3,
Distribution used for $|\mathcal{L}|=27$
("A", "D", "G"): 8, ("A", "D", "H"): 4, ("A", "D", "I"): 2, ("A", "E", "G"): 3, ("A", "E", "H"): 4, ("A", "E", "I"): 5, ("A", "F", "G"): 6, ("A", "F", "H"): 3, ("A", "F", "I"): 6, ("B", "D", "G"): 10, ("B", "D", "H"): 6, ("B", "D", "I"): 1, ("B", "E", "G"): 7, ("B", "E", "H"): 2, ("B", "E", "I"): 4, ("B", "F", "G"): 6, ("B", "F", "H"): 6, ("B", "F", "I"): 3, ("C", "D", "G"): 4, ("C", "D", "H"): 5, ("C", "D", "I"): 6, ("C", "E", "G"): 7, ("C", "E", "H"): 4, ("C", "E", "I"): 2, ("C", "F", "G"): 3, ("C", "F", "H"): 8, ("C", "F", "I"): 5 ,
Distribution used for $|\mathcal{L}|=36$
("A", "E", "H"): 6, ("A", "E", "I"): 8, ("A", "E", "J"): 13, ("A", "F", "H"): 9, ("A", "F", "I"): 1, ("A", "F", "J"): 2, ("A", "G", "H"): 11, ("A", "G", "I"): 2, ("A", "G", "J"): 1, ("B", "E", "H"): 15, ("B", "E", "I"): 4, ("B", "E", "J"): 8, ("B", "F", "H"): 1, ("B", "F", "I"): 5, ("B", "F", "J"): 5, ("B", "G", "H"): 1, ("B", "G", "I"): 7, ("B", "G", "J"): 2, ("C", "E", "H"): 4, ("C", "E", "I"): 5, ("C", "E", "J"): 12, ("C", "F", "H"): 3, ("C", "F", "I"): 2, ("C", "F", "J"): 2, ("C", "G", "H"): 4, ("C", "G", "I"): 8, ("C", "G", "J"): 12, ("D", "E", "H"): 2, ("D", "E", "I"): 4, ("D", "E", "J"): 2, ("D", "F", "H"): 1, ("D", "F", "I"): 2, ("D", "F", "J"): 2, ("D", "G", "H"): 9, ("D", "G", "I"): 2, ("D", "G", "J"): 3
Distribution used for $|\mathcal{L}|=54$
(’A’, ’D’, ’E’, ’H’): 1, (’A’, ’D’, ’E’, ’I’): 5, (’A’, ’D’, ’E’, ’J’): 9, (’A’, ’D’, ’F’, ’H’): 1, (’A’, ’D’, ’F’, ’I’): 10, (’A’, ’D’, ’F’, ’J’): 2, (’A’, ’D’, ’G’, ’H’): 3, (’A’, ’D’, ’G’, ’I’): 2, (’A’, ’D’, ’G’, ’J’): 5, (’A’, ’E’, ’E’, ’H’): 1, (’A’, ’E’, ’E’, ’I’): 1, (’A’, ’E’, ’E’, ’J’): 5, (’A’, ’E’, ’F’, ’H’): 5, (’A’, ’E’, ’F’, ’I’): 12, (’A’, ’E’, ’F’, ’J’): 2, (’A’, ’E’, ’G’, ’H’): 6, (’A’, ’E’, ’G’, ’I’): 4, (’A’, ’E’, ’G’, ’J’): 2, (’B’, ’D’, ’E’, ’H’): 10, (’B’, ’D’, ’E’, ’I’): 2, (’B’, ’D’, ’E’, ’J’): 1, (’B’, ’D’, ’F’, ’H’): 7, (’B’, ’D’, ’F’, ’I’): 2, (’B’, ’D’, ’F’, ’J’): 3, (’B’, ’D’, ’G’, ’H’): 8, (’B’, ’D’, ’G’, ’I’): 2, (’B’, ’D’, ’G’, ’J’): 7, (’B’, ’E’, ’E’, ’H’): 2, (’B’, ’E’, ’E’, ’I’): 2, (’B’, ’E’, ’E’, ’J’): 3, (’B’, ’E’, ’F’, ’H’): 4, (’B’, ’E’, ’F’, ’I’): 5, (’B’, ’E’, ’F’, ’J’): 13, (’B’, ’E’, ’G’, ’H’): 2, (’B’, ’E’, ’G’, ’I’): 4, (’B’, ’E’, ’G’, ’J’): 12, (’C’, ’D’, ’E’, ’H’): 12, (’C’, ’D’, ’E’, ’I’): 1, (’C’, ’D’, ’E’, ’J’): 3, (’C’, ’D’, ’F’, ’H’): 1, (’C’, ’D’, ’F’, ’I’): 15, (’C’, ’D’, ’F’, ’J’): 9, (’C’, ’D’, ’G’, ’H’): 11, (’C’, ’D’, ’G’, ’I’): 1, (’C’, ’D’, ’G’, ’J’): 4, (’C’, ’E’, ’E’, ’H’): 1, (’C’, ’E’, ’E’, ’I’): 5, (’C’, ’E’, ’E’, ’J’): 3, (’C’, ’E’, ’F’, ’H’): 11, (’C’, ’E’, ’F’, ’I’): 13, (’C’, ’E’, ’F’, ’J’): 3, (’C’, ’E’, ’G’, ’H’): 7, (’C’, ’E’, ’G’, ’I’): 4, (’C’, ’E’, ’G’, ’J’): 1
Appendix C LLM Judge Prompt
We used the LLM judge only for the mode identification task, where model responses were often difficult to evaluate using rule-based methods. Many responses implied an answer but didn’t state it clearly or follow a consistent format, making extraction with regular expressions unreliable. For the other two tasks—MLE and sampling—models typically followed the expected output format, or when they didn’t, their responses clearly lacked an answer, making them easy to discard.
To ensure the judge’s reliability, we compared its decisions with human judgments on 100 randomly selected responses. The judge matched human decisions in 99 cases, indicating its reliable judgments.
To automate evaluation, the judge model receives the task, expected output, and model response, and decides whether the response is correct.
Below, is the system prompt used for our judge LLM, Llama3.3-70B
You are an expert evaluator assessing the correctness of responses. Your task is to judge whether a response is correct based on an expected answer. If the response correctly conveys the intended meaning, mark it Correct. Otherwise, mark it Incorrect.
And here’s the template of the user prompt that is given to the judge model:
Given the following question, expected answer, and response, judge whether the response is correct. Clearly state your judgment as "Judgment: Correct." or "Judgment: Incorrect.".
Question: [question]
Expected Answer: [expected]
Response to Evaluate: [response]
Judgment (Correct/Incorrect):
Appendix D Independence of Generated Samples
Table 3: Independence analysis of generated samples by each model. Reported values include the Durbin–Watson (DW) statistic and transition matrix entropy. Samples are classified as non-independent if DW < 1.5 or > 2.5, or if the entropy ratio < 0.8.
Model (DW) statistic entropy-ratio Is independent?
Llama-3.1-8B-Instruct 0.118 0.245 No
Qwen2.5-7B-Instruct-1M 0.002 0.209 No
DeepSeek-R1-Distill-Qwen-7B 0.002 0.195 No
Llama3.3-70B 0.153 0.323 No
GPT-4o-mini 0.636 0.608 No
GPT-4.1-mini 0.149 0.336 No
In our observations, LLM-generated samples are not fully independent and often follow ordering patterns similar to those in the input prompt. For example, as illustrated below, DeepSeek-R1-Distill-Qwen explicitly acknowledges the need to shuffle the samples. However, due to the lack of access to a code interpreter or true randomness, it prioritizes matching target frequencies over enforcing independence:
"So, the final step is to create a list of 84 triplets, where each triplet is repeated according to its count, and then the list is shuffled to randomize the order. But since I can’t perform the shuffling here, I’ll have to present the triplet counts and note that the actual output would be a shuffled list"
To investigate this phenomenon more systematically, we conducted additional analyses using two statistical methods: (1) the Durbin–Watson (DW) statistic [Durbin and Watson, 1951] to detect autocorrelation, and (2) transition matrix entropy [Shannon, 1948] to evaluate diversity in transition patterns. Samples are classified as non-independent based on the aggregated results of these two measures. Table 3 reports the corresponding statistics for samples generated by each model.
Appendix E Evaluation on Real-world Data
Table 4: Frequencies of joint distribution outcomes for the selected subset of the Mushroom dataset. Each column corresponds to one outcome of the joint distribution, along with its frequency in 80 samples. The distribution is highly skewed, with several outcomes having zero frequency.
Mushroom Feature Labels
gill-attachment f f f f f f f f a a a a a a a a
gill-size b b b b n n n n b b b b n n n n
gill-color k g p o k g p o k g p o k g p o
Frequency 13 17 34 0 2 0 9 0 0 0 0 0 0 0 0 0
Table 5: Performance of different models across all six task configurations on the Mushroom dataset. Mode identification tasks are evaluated using accuracy, while MLE and sampling tasks are evaluated using total variation distance (TVD).
Model joint-mode cond-mode joint-MLE cond-MLE joint-samp cond-samp
Llama-3.1-8B-Instruct 80 78 0.011 0.111 0.465 0.414
Qwen2.5-7B-Instruct-1M 95 75 0.003 0.131 0.544 0.441
DeepSeek-R1-Distill-Qwen-7B 96 96 0.002 0.100 0.237 0.173
Llama3.3-70B 100 100 0.0 0.005 0.111 0.089
GPT-4o-mini 100 91 0.0 0.028 0.287 0.352
GPT-4.1-mini 100 100 0.0 0.0 0.193 0.218
To validate our findings on real-world data, we conducted experiments on the Mushroom dataset [mus, 1981], focusing on three categorical features and a subset of 80 samples. The resulting joint distribution is highly skewed, with a large proportion of zero-frequency outcomes, making the task particularly challenging. Table 4 summarizes the observed outcomes and their frequencies in this subset. The experimental results, detailed in Table 5, show that larger models, as well as distilled variants, consistently outperform smaller models across all task configurations.
Appendix F Effect of Sample Count in Context
To investigate the impact of input length independently from task complexity, we fixed the joint space size at $|\mathcal{L}|=12$ and varied the number of observed samples $K$ provided in the prompt. As shown in Figure 4, increasing the number of in-context samples leads to a clear drop in accuracy for both models. This degradation suggests that the challenge stems not just from the probabilistic reasoning task itself, but also from the models’ limited capacity to process and reason over longer contexts. This highlights a core limitation in current LLMs and points to context scaling as a key direction for future improvement.
Figure 4: Accuracy on joint-mode task with fixed joint space size $|\mathcal{L}|=12$ as the number of observed in-context samples increases.
Appendix G Access to Code Interpreter
Table 6: Accuracy of GPT-4.1-mini on the joint-mode task under three configurations across different joint sizes. Results show that accuracy in the samples-in-context setting with access to code interpreter closely matches the frequencies-provided case.
Setting $|\mathcal{L}|=12$ $|\mathcal{L}|=27$ $|\mathcal{L}|=36$ $|\mathcal{L}|=54$
Frequencies in context 1.0 1.0 0.99 0.96
Raw samples in context 0.87 0.71 0.55 0.28
Raw samples + code interpreter 1.0 1.0 0.99 0.97
In the frequency-based setting, both Llama3.3-70B and GPT-4.1-mini achieved near-perfect accuracy, but their performance degraded sharply when asked to infer frequencies directly from raw samples due to limited counting abilities. To explore mitigation strategies, we conducted an additional set of experiments with GPT-4.1-mini augmented with access to a code interpreter. In this setup, prompts encouraged the model to write Python code to count outcome frequencies before solving the joint-mode task. This hybrid approach led to a substantial performance gain. As shown in Table 6, accuracy levels were restored to nearly those observed in the frequency-provided condition, demonstrating that the main limitation lies in the models’ ability to perform reliable counting over long contexts rather than in reasoning about the task itself.